# Planet formation regulated by galactic-scale interstellar turbulence

### v1

@@ 0.17 Sensitivity to the initial disc mass:
We show the resultant disc properties in  Extended Data Figure 5, which again are in good agreement with the observed disc properties. Our adopted initial disc masses are approximately the maximum we could adopt while still remaining consistent with the observed distribution of disc masses. The outcome for the half-life disc replenishment factor is shown in Extended Data Figure 6. Although in this case $f_{\mathrm{M,1/2}}$ is somewhat reduced early on in the disc evolution, overall differences in the distribution over time is minor. The fraction of discs with at least some recently accreted material remains similar, while we find that $f_{\mathrm{M,1/2}}\sim 0.2{-}0.45$ throughout the disc lifetime. Therefore, even under the assumption that the primordial disc is coincidentally similar in mass to the subsequently accreted one, and ignoring all internal depletion processes, we still expect a substantial fraction of discs to be composed of recently captured disc material. Our conclusion that the present day disc properties are not a good probe of internal disc physics therefore appears robust.
@@ 0.18 Disc turbulence and angular momentum transport:
In our model, we have assumed that disc mass is accreted onto the star with a fixed time-scale $\tau_{\mathrm{acc}}$ via an undefined angular momentum transport process. This process could be internal to the star-disc system, driven by magnetic winds or due to (magneto-)hydrodynamic instability.[90, 67] However, the ISM itself could also be the origin of protoplanetary disc turbulence. The degree to which angular momentum transport is mediated by turbulence remains an open question, but measurements indicate it could be the origin of stellar accretion in at least some discs.[7]
To estimate the degree of ISM-induced turbulence, we must compute the balance between the rate at which energy is deposited compared to the decay of that turbulent energy. The rate at which energy decays is:[37]
@@
Then, if the fraction of the energy that goes into turbulent motion in the disc is $\epsilon_{\mathrm{t}}=|\dot{E}_{\mathrm{decay}}/\dot{E}_{\mathrm{in}}|\leq 1$ and the disc is in a steady state, then (rewriting equation 30 of Klessen & Henebelle 2010[37]):
@@
Therefore, the BHL accretion rate $\dot{M}_{\mathrm{BHL}}$ needs to remain sufficient to drive observed turbulence throughout the disc lifetime. However, this requirement is probably met for most discs in the context of our model. With the exception of DM Tau, which has turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, and IM Lup which has $v_{\rm{t,disc}}\sim 0.4{-}0.6\,c_{\mathrm{s}}$,[60] attempts to measure turbulence in protoplanetary discs have resulted in upper limits. For example, for the disc around HD 163296 this upper limit is $v_{\rm{t,disc}}\lesssim 0.05\,c_{\mathrm{s}}$,[114, 56] and for TW Hydra it is $v_{\rm{t,disc}}\lesssim 0.08\,c_{\mathrm{s}}$.[54, 115] For $M_{\mathrm{disc}}/m_{*}=0.01$, $H_{\mathrm{disc}}/R_{\mathrm{disc}}=0.1$ and $\epsilon_{\rm{t}}=1$, then even substituting a modest $\dot{M}_{\rm BHL}\sim 10^{-11}\,M_{\odot}$ yr-1 yields $v_{\mathrm{t,disc}}\sim 0.03c_{\mathrm{s}}$, similar to the empirical upper limit on turbulence in some discs. This is approximately the median BHL accretion rate for a solar mass star in our model at an age of $\sim 3$ Myr (see the bottom panel of Figure 1). Typically values of $M_{\mathrm{disc}}$ are even smaller than $10^{-2}m_{*}$, so even greater turbulence may be imparted by the ISM. Broadly, for discs older than $\sim 3$ Myr (as is the case for all those with upper limits on the turbulence) we expect relatively low levels of turbulence, while for younger discs (like DM Tau[59] and IM Lup[61]) turbulence driven by the ISM can also be larger.
@@
although in practice $|\dot{E}_{\mathrm{acc}}|\ll|\dot{E}_{\mathrm{decay}}|$. Then we can solve the initial value problem, including the mass evolution of the disc as before. In principle, we could then compute the disc evolution fully consistently assuming only the ISM as the driver of internal physics, since we then have $\alpha_{\mathrm{SS}}$ or equivalently $\tau_{\mathrm{acc}}$. However, in practice, this requires introducing a free parameter $\epsilon_{\mathrm{t}}$, which may not be a constant. Strictly, it also means self-consistently evolving the disc outer radius (or disc scale radius), which should viscously expand, such that we can compute the correct $v_{\mathrm{in}}$. This would also be balanced by the contraction due to deposition of low angular momentum material,[66] resulting in additional (non-turbulent) stellar accretion. For these reasons, self-consistently solving the disc evolution would considerably complicate our model. We therefore leave this to future work. Nonetheless, we can adopt sensible parameters: $\epsilon_{\mathrm{t}}=1$, $R_{\mathrm{disc}}=250(m_{*}/1\,M_{\odot})$ au and $H_{\mathrm{disc}}=0.1R_{\mathrm{disc}}$ with our fixed $\tau_{\mathrm{acc}}$ to obtain an estimate for the maximal turbulent disc velocity $v_{\mathrm{t,disc}}$ over time. The outcome of this exercise is shown in Figure 4, and discussed in the main article. We find that sufficient turbulent energy is supplied to discs from the ISM to sustain the observed turbulence.
We can also consider the specific case of DM Tau, which is a star of mass $\sim 0.4{-}0.5\,M_{\odot}$ hosting a disc with a dust cavity and accreting at a rate $\sim 10^{-8}\,M_{\odot}$ yr-1.[116, 117] It has a turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, or viscous $\alpha_{\mathrm{SS}}\sim 0.1$.[6] For this disc, a recent and relatively rapid period of accretion would be needed to drive the observed turbulence. Estimating $M_{\mathrm{disc}}/m_{*}\approx 0.03$,[118] this would imply recent accretion from the ISM at a rate $\dot{M}_{\mathrm{BHL}}\gtrsim 2\times 10^{-8}\,M_{\odot}$ yr-1 that ended in the last $\sim 10^{4}$ years. This is similar to the current stellar accretion rate,[116] suggesting that the majority of the mass of the present day disc could have originated from this putative in-fall event. This scenario is plausible statistically within the context of our model. However, apart from the argument above, there is currently no particular evidence that DM Tau has undergone recent in-fall. This underlines the importance of a systematic search for observational signatures of late stage in-fall events.
@@ 0.19 Observability of in-fall events:
A valid question to ask is the degree to which the model we present in this work predicts more abundant evidence of late-stage in-fall events from the data that is currently available. In fact, growing evidence suggests that a considerable fraction of discs are undergoing some kind of interaction with their environment. For example, a recent scattered light survey of NIR polarimetric imaging for 43 discs in Taurus reported that $16$ percent of those discs showed evidence of interaction with ambient material (although this sample is not necessarily representative).[32] While inferring observational signatures of late stage in-fall requires full hydrodynamic and radiative transfer simulations,[119] we can make a simple estimate for the density of gas we expect to be surrounding discs in our model. For accretion onto a disc from a radius $R_{\mathrm{acc}}$, the maximum free-fall time onto a star of mass $m_{*}$ is:
@@
The typical column density implied by equation 23 is relatively low, such that we would expect that in-fall is often challenging to trace observationally. We may consider the example of CO isotopologues, which are often the brightest tracers used to probe the gas content of protoplanetary discs with ALMA. These isotopologues have a self-shielding column density of $\sim 10^{15}$ cm-2,[120, 121] and typical abundance relative to molecular hydrogen $\sim 10^{-5}$.[104] As an order of magnitude estimate, we may therefore expect column densities $\lesssim 10^{20}$ cm-2 to have the CO photodissociated. From equation 23, we plot the $N_{\mathrm{ext}}$ distribution over time resulting from the accretion rates in our model in Extended Data Figure 7. We see immediately in-falling CO would be photodissociated for the majority of discs in the age range $\sim 1{-}3$ Myr. Interestingly, in this age range, the fraction of discs above the threshold $N_{\mathrm{ext}}=10^{20}$ cm-2, approximately the CO self-shielding column, drops from approximately $30$ percent to close to 00 percent. Given that this is the range of ages in Taurus, this would be commensurate with the inferred $\sim 16$ percent of discs with evidence of interaction with the ambient medium from scattered light observations.[32] However, the physical column density down to which these observations are sensitive is not clear. While these estimates are approximate, they highlight the importance of deep surveys of the environments surrounding large numbers of discs. Such data could be applied to make statistical inferences about the prevalence of late stage in-fall, testing the theoretical predictions we have made in this work.
@@ 0.20 Caveats for the ISM accretion rate:
BHL accretion mediated by a disc: The nature of BHL accretion has a long history of testing with numerical experiments. In a supersonic, uniform medium, these broadly validate the analytic estimate for the BHL accretion rate within a factor of order unity.[122, 123, 124, 125, 34, 126] In a turbulent medium, the instantaneous accretion rate may be limited by the local vorticity instead of the velocity.[43] In this case there may be a substantial reduction in the accretion rate, but only when the Mach number is low (i.e. $\Delta v\sim c_{\mathrm{s}}$), which is rarely the case in our numerical calculations. In a magnetised medium, the BHL accretion rate may also be reduced by a factor $\sim 2$, although this is dependent on the ratio of magnetic to thermal pressure.[75, 76] Overall, we expect our simple estimates to yield a reasonable approximation for the mass accretion rate.
A further pertinent question is what we expect the appearance of the resultant disc to be. As previously discussed, when accreting material with with a high vorticity, the resultant disc outer radius can be of order the BHL radius – see, for example, Figure 3 of Krumholz et al. (2006).[43] What is more uncertain is what should be the structure of a disc if the angular momentum of the accreted material is not aligned with the existing material. Naively, such randomised angular momentum could result in ‘puffed up’ discs. By contrast, at least the mm-dust component of discs is frequently observed to be extremely flat.[127, 128, 129, 5, 62] However, mm-dust is also commonly more centrally concentrated than the gaseous disc by a factor $\sim 2.5$.[130, 131, 48] Since dust radial drift, growth and vertical settling should occur simultaneously, it is not clear how accretion of turbulent gas in the outer disc should be reflected in the dust distribution. This represents an important topic for future study.
Cloudlet capture: Another possible avenue for the capture of interstellar gas is the capture of small cloudlets via tidal energy dissipation. During the close passage between a star and a cloud, the orbital energy may be dissipated by non-radial oscillations within the cloud, analogous to the formation of tight binaries in globular clusters,[132, 133] in a process that remains similar across a range of internal density profiles.[134] Capture requires a periastron $a_{\mathrm{p}}$ during an encounter between a cloud with radius $R_{\mathrm{cloud}}$ and mass $M_{\mathrm{cloud}}$ and star with mass $m_{*}$ such that:

### v2

@@ 3.3 Replenished mass fraction:
Figure 7: The fraction of mass that has been accreted in the last half of the disc life-time, $f_{M,1/2}$, as a function of time for $500$ discs in our model. On the left, we show the time evolution of $f_{M,1/2}$ for a random subset of $100$ discs for clarity. These lines are coloured by the ratio of the disc mass to stellar mass $M_{\mathrm{disc}}/m_{*}$, shown in the colour bar along the top. On the right hand side we show a cumulative fraction for $f_{M,1/2}$ at $1$, $2$, $3$ and $5$ Myr. Only discs with a total mass $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$ are included in this cumulative fraction, lower mass discs are assumed to be dispersed.
@@ 3.4 Sensitivity to the initial disc mass:
We show the resultant disc properties in  8, which again are in good agreement with the observed disc properties. Our adopted initial disc masses are approximately the maximum we could adopt while still remaining consistent with the observed distribution of disc masses. The outcome for the half-life disc replenishment factor is shown in Figure 9. Although in this case $f_{\mathrm{M,1/2}}$ is somewhat reduced early on in the disc evolution, overall differences in the distribution over time is minor. The fraction of discs with at least some recently accreted material remains similar, while we find that $f_{\mathrm{M,1/2}}\sim 0.2{-}0.45$ throughout the disc lifetime. Therefore, even under the assumption that the primordial disc is coincidentally similar in mass to the subsequently accreted one, and ignoring all internal depletion processes, we still expect a substantial fraction of discs to be composed of recently captured disc material. Our conclusion that the present day disc properties are not a good probe of internal disc physics therefore appears robust.
Figure 8: As in Figure 3, but for an initial disc mass $M_{\mathrm{disc}}(t=0)=0.01(m_{*}/1\,M_{\odot})^{2}\,M_{\odot}$ with $1$ dex of scatter.
Figure 9: As in Figure 7, but for an initial disc mass $M_{\mathrm{disc}}(t=0)=0.01(m_{*}/1\,M_{\odot})^{2}\,M_{\odot}$ with $1$ dex of scatter.
Figure 10: The column number density of material assuming a mean molecular mass $\mu=2.3$ and $\epsilon_{\mathrm{geo}}=1$ following equation 21 for the surviving discs in our fiducial model. The colours of the lines represent the age of the discs. The red vertical dashed line represents an approximate column density for photodissociation of 12CO, based on a self-shielding column of $10^{15}$ cm-2 (van Dishoeck & Black, 1988) and abundance relative to hydrogen of $10^{-5}$ (Trapman et al., 2023)
@@ 3.5 Disc turbulence and angular momentum transport:
In our model, we have assumed that disc mass is accreted onto the star with a fixed time-scale $\tau_{\mathrm{acc}}$ via an undefined angular momentum transport process. This process could be internal to the star-disc system, driven by magnetic winds or due to (magneto-)hydrodynamic instability (Tabone et al., 2022; Lesur et al., 2023). However, the ISM itself could also be the origin of protoplanetary disc turbulence. The degree to which angular momentum transport is mediated by turbulence remains an open question, but measurements indicate it could be the origin of stellar accretion in at least some discs (Rosotti, 2023).
To estimate the degree of ISM-induced turbulence, we must compute the balance between the rate at which energy is deposited compared to the decay of that turbulent energy. The rate at which energy decays is (Klessen & Hennebelle, 2010):
@@
Then, if the fraction of the energy that goes into turbulent motion in the disc is $\epsilon_{\mathrm{t}}=|\dot{E}_{\mathrm{decay}}/\dot{E}_{\mathrm{in}}|\leq 1$ and the disc is in a steady state, then (rewriting equation 30 of Klessen & Hennebelle 2010):
@@
Therefore, the BHL accretion rate $\dot{M}_{\mathrm{BHL}}$ needs to remain sufficient to drive observed turbulence throughout the disc lifetime. However, this requirement is probably met for most discs in the context of our model. With the exception of DM Tau, which has turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, and IM Lup which has $v_{\rm{t,disc}}\sim 0.4{-}0.6\,c_{\mathrm{s}}$ (Paneque-Carreño et al., 2024) attempts to measure turbulence in protoplanetary discs have resulted in upper limits. For example, for the disc around HD 163296 this upper limit is $v_{\rm{t,disc}}\lesssim 0.05\,c_{\mathrm{s}}$ (Flaherty et al., 2015, 2017) and for TW Hydra it is $v_{\rm{t,disc}}\lesssim 0.08\,c_{\mathrm{s}}$ (Flaherty et al., 2018; Teague et al., 2018). For $M_{\mathrm{disc}}/m_{*}=0.01$, $H_{\mathrm{disc}}/R_{\mathrm{disc}}=0.1$ and $\epsilon_{\rm{t}}=1$, then even substituting a modest $\dot{M}_{\rm BHL}\sim 10^{-11}\,M_{\odot}$ yr-1 yields $v_{\mathrm{t,disc}}\sim 0.03c_{\mathrm{s}}$, similar to the empirical upper limit on turbulence in some discs. This is approximately the median BHL accretion rate for a solar mass star in our model at an age of $\sim 3$ Myr (see the bottom panel of Figure 2). Typically values of $M_{\mathrm{disc}}$ are even smaller than $10^{-2}m_{*}$, so even greater turbulence may be imparted by the ISM. Broadly, for discs older than $\sim 3$ Myr (as is the case for all those with upper limits on the turbulence) we expect relatively low levels of turbulence, while for younger discs (like DM Tau (Zhang et al., 2020) and IM Lup (Mawet et al., 2012)) turbulence driven by the ISM can also be larger.
@@
although in practice $|\dot{E}_{\mathrm{acc}}|\ll|\dot{E}_{\mathrm{decay}}|$. Then we can solve the initial value problem, including the mass evolution of the disc as before. In principle, we could then compute the disc evolution fully consistently assuming only the ISM as the driver of internal physics, since we then have $\alpha_{\mathrm{SS}}$ or equivalently $\tau_{\mathrm{acc}}$. However, in practice, this requires introducing a free parameter $\epsilon_{\mathrm{t}}$, which may not be a constant. Strictly, it also means self-consistently evolving the disc outer radius (or disc scale radius), which should viscously expand, such that we can compute the correct $v_{\mathrm{in}}$. This would also be balanced by the contraction due to deposition of low angular momentum material (Wijnen et al., 2017), resulting in additional (non-turbulent) stellar accretion. For these reasons, self-consistently solving the disc evolution would considerably complicate our model. We therefore leave this to future work. Nonetheless, we can adopt sensible parameters: $\epsilon_{\mathrm{t}}=1$, $R_{\mathrm{disc}}=250(m_{*}/1\,M_{\odot})$ au and $H_{\mathrm{disc}}=0.1R_{\mathrm{disc}}$ with our fixed $\tau_{\mathrm{acc}}$ to obtain an estimate for the maximal turbulent disc velocity $v_{\mathrm{t,disc}}$ over time. The outcome of this exercise is shown in Figure 11, and discussed in the main article. We find that sufficient turbulent energy is supplied to discs from the ISM to sustain the observed turbulence.
Figure 11: Evolution of the cumulative distribution function for the turbulent Mach number (or viscous $\alpha_{\mathrm{SS}}$, top axis) for the discs in our model, assuming a maximal energy input efficiency $\epsilon_{\mathrm{t}}=1$. Each line colour represents a different disc age in our model, from $0.1$ to $9$ Myr. We show as coloured vertical lines some examples of discs with constraints on their turbulent velocity from the literature, several of which are upper limits (Rosotti, 2023).
We can also consider the specific case of DM Tau, which is a star of mass $\sim 0.4{-}0.5\,M_{\odot}$ hosting a disc with a dust cavity and accreting at a rate $\sim 10^{-8}\,M_{\odot}$ yr-1 (Manara et al., 2014; Francis & van der Marel, 2020). It has a turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, or viscous $\alpha_{\mathrm{SS}}\sim 0.1$ (Flaherty et al., 2020). For this disc, a recent and relatively rapid period of accretion would be needed to drive the observed turbulence. Estimating $M_{\mathrm{disc}}/m_{*}\approx 0.03$ (Andrews et al., 2013) this would imply recent accretion from the ISM at a rate $\dot{M}_{\mathrm{BHL}}\gtrsim 2\times 10^{-8}\,M_{\odot}$ yr-1 that ended in the last $\sim 10^{4}$ years. This is similar to the current stellar accretion rate (Manara et al., 2014), suggesting that the majority of the mass of the present day disc could have originated from this putative in-fall event. This scenario is plausible statistically within the context of our model. However, apart from the argument above, there is currently no particular evidence that DM Tau has undergone recent in-fall. This underlines the importance of a systematic search for observational signatures of late stage in-fall events.
@@ 3.6 Observability of in-fall events:
A valid question to ask is the degree to which the model we present in this work predicts more abundant evidence of late-stage in-fall events from the data that is currently available. In fact, growing evidence suggests that a considerable fraction of discs are undergoing some kind of interaction with their environment. For example, from their recent scattered light survey of NIR polarimetric imaging for 43 discs in Taurus, Garufi et al. (2024) reported that $16$ percent of those discs showed evidence of interaction with ambient material (although this sample is not necessarily representative). While inferring observational signatures of late stage in-fall requires full hydrodynamic and radiative transfer simulations (Krieger et al., 2024), we can make a simple estimate for the density of gas we expect to be surrounding discs in our model. For accretion onto a disc from a radius $R_{\mathrm{acc}}$, the maximum free-fall time onto a star of mass $m_{*}$ is:
@@
Figure 12: The column number density of material assuming a mean molecular mass $\mu=2.3$ and $\epsilon_{\mathrm{geo}}=1$ following equation 21 for the surviving discs in our fiducial model. The colours of the lines represent the age of the discs. The red vertical dashed line represents an approximate column density for photodissociation of 12CO, based on a self-shielding column of $10^{15}$ cm-2 (van Dishoeck & Black, 1988) and abundance relative to hydrogen of $10^{-5}$ (Trapman et al., 2023).
The typical column density implied by equation 21 is relatively low, such that we would expect that in-fall is often challenging to trace observationally. We may consider the example of CO isotopologues, which are often the brightest tracers used to probe the gas content of protoplanetary discs with ALMA. These isotopologues have a self-shielding column density of $\sim 10^{15}$ cm-2 (van Dishoeck & Black, 1988; Visser et al., 2009), and typical abundance relative to molecular hydrogen $\sim 10^{-5}$ (Trapman et al., 2023). As an order of magnitude estimate, we may therefore expect column densities $\lesssim 10^{20}$ cm-2 to have the CO photodissociated. From equation 21, we plot the $N_{\mathrm{ext}}$ distribution over time resulting from the accretion rates in our model in Figure 12. We see immediately in-falling CO would be photodissociated for the majority of discs in the age range $\sim 1{-}3$ Myr. Interestingly, in this age range, the fraction of discs above the threshold $N_{\mathrm{ext}}=10^{20}$ cm-2, approximately the CO self-shielding column, drops from approximately $30$ percent to close to 00 percent. Given that this is the range of ages in Taurus, this would be commensurate with the inferred $\sim 16$ percent of discs with evidence of interaction with the ambient medium from scattered light observations (Garufi et al., 2024). However, the physical column density down to which these observations are sensitive is not clear. While these estimates are approximate, they highlight the importance of deep surveys of the environments surrounding large numbers of discs. Such data could be applied to make statistical inferences about the prevalence of late stage in-fall, testing the theoretical predictions we have made in this work.
4 Discussion
@@ 4.1 Caveats for the ISM accretion rate:
4.1.1 BHL accretion mediated by a disc
The nature of BHL accretion has a long history of testing with numerical experiments. In a supersonic, uniform medium, these broadly validate the analytic estimate for the BHL accretion rate within a factor of order unity (Hunt, 1971, 1979; Ruffert, 1994, 1996; Edgar, 2004; El Mellah & Casse, 2015). In a turbulent medium, the instantaneous accretion rate may be limited by the local vorticity instead of the velocity (Krumholz et al., 2006). In this case there may be a substantial reduction in the accretion rate, but only when the Mach number is low (i.e. $\Delta v\sim c_{\mathrm{s}}$), which is rarely the case in our numerical calculations. In a magnetised medium, the BHL accretion rate may also be reduced by a factor $\sim 2$, although this is dependent on the ratio of magnetic to thermal pressure (Lee et al., 2014; Burleigh et al., 2017). Overall, we expect our simple estimates to yield a reasonable approximation for the mass accretion rate.
A further pertinent question is what we expect the appearance of the resultant disc to be. As previously discussed, when accreting material with with a high vorticity, the resultant disc outer radius can be of order the BHL radius (Krumholz et al., 2006; Kuffmeier et al., 2023). What is more uncertain is what should be the structure of a disc if the angular momentum of the accreted material is not aligned with the existing material. Depending on the relative angular momentum vector, this could result in ‘puffed up’ discs, or misalignment between the inner and outer regions (Kuffmeier et al., 2021). By contrast, at least the mm-dust component of discs is frequently observed to be extremely flat (Duchêne et al., 2003; Pinte et al., 2007, 2016; Villenave et al., 2020, 2022). However, mm-dust is also commonly more centrally concentrated than the gaseous disc by a factor $\sim 2.5$ (Ansdell et al., 2018; Facchini et al., 2019; Andrews, 2020). Since dust radial drift, growth and vertical settling should occur simultaneously, it is not clear how accretion of turbulent gas in the outer disc should be reflected in the dust distribution. This represents an important topic for future study.
4.1.2 Cloudlet capture
Another possible avenue for the capture of interstellar gas is the capture of small cloudlets via tidal energy dissipation. During the close passage between a star and a cloud, the orbital energy may be dissipated by non-radial oscillations within the cloud, analogous to the formation of tight binaries in globular clusters (Robe, 1968; Fabian et al., 1975), in a process that remains similar across a range of internal density profiles (Lee & Ostriker, 1986). Capture requires a periastron $a_{\mathrm{p}}$ during an encounter between a cloud with radius $R_{\mathrm{cloud}}$ and mass $M_{\mathrm{cloud}}$ and star with mass $m_{*}$ such that: